leila123456:

A six-sided number cube is rolled twice. Which expression can be used to find P(5, then3)?

justjm:

I believe it is asking you to find \(\text{nPr(5,3)}\) Just use the formula for permutations \[\text{nPr(5,3)}=\frac{5!}{(5-3)!}\] \[=\frac{5 \times 4 \times 3 \times \cancel{2!}}{\cancel{2!}}\] \[=5 \times 4 \times 3\] You can evaluate that